# Fourier Learning Machines: Nonharmonic Fourier-Based Neural Networks for Scientific Machine Learning

### v1

@@ V Tanh Vanilla feedforward NN with Tanh activation function and 3 hidden layers
When numerically solving each PDE, we trained FLMs and other NN models for 18–27 different hyperparameter configurations, with a maximum of 10,000 epochs and a training loss tolerance of $10^{-4}$. All error metrics (MSE, MAE, Max Error) and the number of epochs required to reach the specified training loss tolerance were averaged over 10 runs with different random seeds. The performances of all tested models applied to each PDE are reported in Tables 5– 20 in Appendix B.
With the best best performing hyperparameter configuration (highlighted in the tables in Appendix B) selected by following the aforementioned procedure, we train the NN models for $30,000$ new additional epochs with a training loss tolerance of $10^{-8}$, and report the error metrics for each PDE in their respective tables.
Example 5.1 (One–Dimensional Heat Equation).
@@
Fig. 4 illustrates the comparison between the exact solution and the FLM approximated solution. The performances of all models are summarized in Table 2.
Figure 4: Solution surface of the Heat equation: (left) exact, (middle) FLM approximation, (right) absolute error.
Table 2: Approximation Errors for the Heat PDE
Models MSE MAE Max Error
FLM $6.24\times 10^{-8}$ $2.01\times 10^{-4}$ $7.37\times 10^{-4}$
SIREN–1 $5.07\times 10^{-7}$ $5.29\times 10^{-4}$ $1.97\times 10^{-3}$
SIREN–2 $3.07\times 10^{-6}$ $1.18\times 10^{-3}$ $8.06\times 10^{-3}$
SIREN–3 $5.40\times 10^{-5}$ $3.90\times 10^{-3}$ $2.00\times 10^{-2}$
V Relu $6.39\times 10^{-2}$ $2.03\times 10^{-1}$ $6.72\times 10^{-1}$
V Tan $9.24\times 10^{-6}$ $1.39\times 10^{-3}$ $4.08\times 10^{-3}$
Example 5.2 (Two-Dimensional Poisson Equation).
@@
Fig. 5 shows the comparison between the exact solution and the FLM approximated solution. The results are summarized in Table 3.
Figure 5: Solution surface of the Poisson equation: (left) exact, (middle) FLM approximation, (right) absolute error.
Table 3: Approximation Errors for the Poisson PDE
Models MSE MAE Max Error
FLM $1.34\times 10^{-7}$ $2.04\times 10^{-4}$ $6.74\times 10^{-4}$
SIREN–1 $2.01\times 10^{-7}$ $3.95\times 10^{-4}$ $1.86\times 10^{-3}$
SIREN–2 $8.07\times 10^{-5}$ $7.09\times 10^{-3}$ $3.43\times 10^{-2}$
V Relu $2.45\times 10^{-1}$ $3.97\times 10^{-1}$ $9.99\times 10^{-1}$
V Tan $2.64\times 10^{-4}$ $9.95\times 10^{-3}$ $6.56\times 10^{-2}$
Example 5.3 (Generalized Black-Scholes PDE).
Consider the PDE extracted form [21], with initial and boundary condition
@@
The comparison between the exact solution and the FLM approximated solution is presented in Fig. 6. The results are presented in Table 4.
Figure 6: Solution surface of the GBS equation: (left) exact, (middle) FLM approximation, (right) absolute error.
Table 4: Approximation Errors for the Generalized Black-Scholes PDE
@@
The numerical experiments demonstrate that FLMs consistently outperform other NN models across a set of multidimensional PDEs. While SIREN and vanilla feedforward NNs perform well when the target function contains similar mathematical behavior as the activation function they use, FLMs achieve significantly lower error metrics and faster convergence (see tables in Appendix B) regardless of the PDE. In terms of the MSE and MAE metrics, FLMs were shown to perform better than all models considered for each PDE. In terms of Max Error, FLMs also showed an overall superior performance, even though not always dominating the other models. In general, out of all of the competing models, SIREN NNs are the models with the closest performance to that of FLMs. In the next section, we apply the proposed FLM model to the solution of a family of OCPs.
5.2 Numerical Experiments for Optimal Control Problem
@@
Both NN and Fourier series have been proven to be very effective to solve OCPs in recent years [31, 22, 23, 24, 25]. To show the effectiveness of FLMs to solve OCPs, we consider the optimal control of a system governed by the dynamics of a 3–strategy odd–circulant evolutionary game, namely, a Rock–Paper–Scissors game [27]. The state of the system, a vector $\boldsymbol{u}(t)=(u_{1},u_{2},u_{3})^{\top}$ with $u_{1}+u_{2}+u_{3}=1$, represents the proportion of a population playing each strategy. The goal is to push these proportions toward the desired equilibrium point $\boldsymbol{u}^{*}=(1/3,1/3,1/3)^{\top}$ representing an equally balanced population.
@@
The system’s evolution is described by the replicator dynamics and is influenced by an external control input, $\gamma(t)$. This control actively alters the game’s payoff matrix, which is defined as $\boldsymbol{A}_{3}(\gamma)=\boldsymbol{L}_{3}+\gamma\boldsymbol{M}_{3}$, where $\boldsymbol{L}_{3}$ is the standard payoff matrix and $\boldsymbol{M}_{3}$ is an actuating matrix and given by
$\boldsymbol{L}_{3}=\begin{pmatrix}0&-1&1\\ 1&0&-1\\ -1&1&0\end{pmatrix},\quad\boldsymbol{M}_{3}=\begin{pmatrix}0&0&1\\ 1&0&0\\ 0&1&0\end{pmatrix}$.
@@
The objective is to find an optimal control strategy $\gamma(t)$ that minimizes a cost functional over a time horizon $[0,T]$. This cost has two components: a term that penalizes the squared distance of the state $\boldsymbol{u}(t)$ from the target equilibrium $\boldsymbol{u}^{*}$, and a term that penalizes the squared control effort $\gamma(t)^{2}$.
@@
An important point to note is that the given OCP possesses a cyclic symmetry. If $\boldsymbol{u}^{*}(t)=(u^{*}_{1}(t),u^{*}_{2}(t),u^{*}_{3}(t))^{\top}$ is the optimal state trajectory, $\gamma^{*}(t)$ is the optimal control and $J^{*}$ is the optimal objective value for an initial condition $\boldsymbol{u_{0}}=(u_{01},u_{02},u_{03})^{\top}$, then any cyclic permutation of this initial state will yield the same optimal control solution and objective value. Consequently, the new optimal state trajectory will be the corresponding cyclic permutation of the original trajectory, $\boldsymbol{u}^{*}(t)$. For instance, if the initial state is permuted to $(u_{02},u_{03},u_{01})$, the optimal trajectory becomes $(u^{*}_{2}(t),u^{*}_{3}(t),u^{*}_{1}(t))$, while the optimal control and objective value remain unchanged. The numerical approximation of the solution of the OCP with our proposed model is presented in the next example.
Example 5.4.
@@
Consider a class of the OCP described above, where the initial condition $\boldsymbol{u_{0}}$ varies withing a given range, written as
min $J(\boldsymbol{u},\gamma)=\int_{0}^{T}\left(\frac{1}{2}\|\boldsymbol{u}-\boldsymbol{u}^{*}\|_{2}^{2}+\frac{r}{2}\gamma^{2}\right)dt$
s.t. $\dot{\boldsymbol{u}}=\boldsymbol{F}(\boldsymbol{u})+\gamma(t)\boldsymbol{G}(\boldsymbol{u})$
$\boldsymbol{u}(0)=\boldsymbol{u}_{0}\in U_{0},$
where $\boldsymbol{u}(t)\in\Delta_{2}$ lies on the 2–simplex. The vector fields, for $i=1,2,3,$ are given by
$\boldsymbol{F}_{i}(\boldsymbol{u})$ $=u_{i}\left((\boldsymbol{e}_{i}-\boldsymbol{u})^{T}\boldsymbol{L}_{3}\boldsymbol{u}\right),$
$\boldsymbol{G}_{i}(\boldsymbol{u})$ $=u_{i}\left((\boldsymbol{e}_{i}-\boldsymbol{u})^{T}\boldsymbol{M}_{3}\boldsymbol{u}\right).$
@@
Let $\widehat{u}_{i}(t,\boldsymbol{u_{0}};\Theta_{u_{i}})$ be the FLM that learns the approximation of state variable $u_{i}(t,u_{0})$, i.e., $\boldsymbol{u}\approx\boldsymbol{\widehat{u}}=\left(\widehat{u}_{1},\widehat{u}_{2},\widehat{u}_{3}\right)^{\top}$ and $\widehat{\gamma}(t,\boldsymbol{u_{0}};\Theta_{\gamma})$ be the FLM that learns the control variable $\gamma(t,u_{0})$ for any trajectory departing from $U_{0}$, where $\Theta=\{\Theta_{u_{1}},\Theta_{u_{2}},\Theta_{u_{3}},\Theta_{\gamma}\}$ is the collection of all network parameters of the respective networks. Given the complex nature of the OCP above, the training process is different than it was for the PDEs and so is the construction of the training loss function. We train all the FLM simultaneously by minimizing the training loss using a penalty method (Ch. 9 in [32]). We define an unconstrained loss function that incorporates the objective functional and penalties for dynamics and initial condition violations as
$\mathcal{L}(\Theta)=\frac{1}{N_{\text{IC}}}\sum_{\boldsymbol{u}_{0}\in U_{0}}\left[\mathcal{J}+\frac{1}{2}\boldsymbol{\mu^{\top}_{1}}\boldsymbol{\mathcal{V}}_{\text{dyn}}+\frac{1}{2}\boldsymbol{\mu^{\top}_{2}}\boldsymbol{\mathcal{V}}_{\text{init}}\right],$ (16)
@@
$\mathcal{J}(\Theta)$ $=\int_{0}^{T}\left(\frac{1}{2}\|\boldsymbol{\widehat{u}}-\boldsymbol{u}^{*}\|^{2}+\frac{r}{2}\widehat{\gamma}^{2}\right)dt,$
$\boldsymbol{\mathcal{V}}_{\text{dyn}}(\Theta)$ $=\int_{0}^{T}\left(\dot{\widehat{\boldsymbol{u}}}-\boldsymbol{F}(\widehat{\boldsymbol{u}})-\widehat{\gamma}\boldsymbol{G}(\widehat{\boldsymbol{u}})\right)\odot\left(\dot{\widehat{\boldsymbol{u}}}-\boldsymbol{F}(\widehat{\boldsymbol{u}})-\widehat{\gamma}\boldsymbol{G}(\widehat{\boldsymbol{u}})\right)dt,$
$\boldsymbol{\mathcal{V}}_{\text{init}}(\Theta)$ $=\left(\boldsymbol{\widehat{u}}(0)-\boldsymbol{u}_{0}\right)\odot\left(\boldsymbol{\widehat{u}}(0)-\boldsymbol{u}_{0}\right),$
where $\boldsymbol{\mu_{1}},\boldsymbol{\mu_{2}}$ are vectors of positive penalty coefficients. Minimizing the total loss $\mathcal{L}$ with respect to $\Theta_{u}$ and $\Theta_{\gamma}$ forces the FLM to approximate the optimal state trajectory and control signal that solve the original OCP. Next, we consider two cases: a fixed initial condition case, for a given $\boldsymbol{u_{0}}$, and a for a range of initial conditions $U_{0}\subset\Delta_{2}$.
@@ 5.2.1 Fixed Initial Condition OCP
First we solve the OCP for fixed initial condition with a one–dimensional FLM, which is a NN with one hidden–layer and cosinecosine activation function. We consider three such FLMs, two for the state variables $u_{1}$ and $u_{2}$ and one for the control $\gamma$. The third state variable is calculated using the relation $u_{3}=1-u_{1}-u_{2}$. The training is performed with default learning rate $0.001$ and betas $\beta_{1}=0.99$ and $\beta_{2}=0.999$ with the ADAM optimizer as well. To test the accuracy of our model, we compare its performance against a numerical solution obtained via PMP to formulate a two–point boundary value problem, which was then solved numerically to a high precision (tolerance of $10^{-8}$) using the solve_bvp function from Python’s SciPy library. The FLM approximated solutions and the PMP–based numerical solutions for three different, fixed initial conditions are shown in Fig. 7. The approximated optimal objective function values have absolute percentage errors of $0.47\%$, $0.37\%$, and $0.34\%$ respectively. Fig. 8 shows the state trajectory for the initial point $\boldsymbol{u}_{0}=(0.20,0.20,0.60)^{\top}$. The parameters of the OCP, as in [27], are chosen as $T=6.0$, $r=0.2$.

### v3

@@ V Tanh Vanilla feedforward NN with Tanh activation function and 3 hidden layers
We trained FLMs and diverse NN models with multiple hyperparameter configurations, selected the best using a hierarchical statistical testing procedure, and further trained the top configuration for 30,000 epochs with a stricter loss tolerance. Details are shown in Appendix C.
Example 5.1 (One–Dimensional Heat Equation)
@@
Fig. 5(a) illustrates the comparison between the exact solution and the FLM approximated solution. The performances of all models are summarized in Table 2(a).
Example 5.2 (Two-Dimensional Poisson Equation)
@@
Fig. 5(b) shows the comparison between the exact solution and the FLM approximated solution. The results are summarized in Table 2(b).
Example 5.3 (Generalized Black-Scholes Equation)
Consider the PDE extracted from Ma et al. (2022), with initial and boundary condition
@@
The comparison between the exact solution and the FLM approximated solution is presented in Fig. 5(c). The results are presented in Table 2(c).
Example 5.4 (Inviscid Burgers’ Equation)
$\frac{\partial u}{\partial t}+u\,\frac{\partial u}{\partial x}=0,\qquad 0\leq x\leq 1,\;0\leq t\leq 1,$
$u(x,0)=1.0+0.35\sin(2\pi x),$
$u(0,t)=u(1,t).$
This PDE has discontinuous solution and the boundary condition $u(0,t)=u(1,t)$ induces a discontinuity in the boundary as well. In this case, we compare the approximated solution with the numerical solution obtained by Godunov’s method. The results are illustrated in Fig. 5(d) and summarized in Table 2(d).
Table 2: Approximation Errors for Various PDEs Using Different Models
(a) Heat PDE
Models MSE MAE Max Error
FLM $6.24\times 10^{-8}$ $2.01\times 10^{-4}$ $7.37\times 10^{-4}$
SIREN–1 $5.07\times 10^{-7}$ $5.29\times 10^{-4}$ $1.97\times 10^{-3}$
SIREN–2 $3.07\times 10^{-6}$ $1.18\times 10^{-3}$ $8.06\times 10^{-3}$
SIREN–3 $5.40\times 10^{-5}$ $3.90\times 10^{-3}$ $2.00\times 10^{-2}$
V Relu $6.39\times 10^{-2}$ $2.03\times 10^{-1}$ $6.72\times 10^{-1}$
V Tan $9.24\times 10^{-6}$ $1.39\times 10^{-3}$ $4.08\times 10^{-3}$
(b) Poisson PDE
Models MSE MAE Max Error
FLM $1.34\times 10^{-7}$ $2.04\times 10^{-4}$ $6.74\times 10^{-4}$
SIREN–1 $2.01\times 10^{-7}$ $3.95\times 10^{-4}$ $1.86\times 10^{-3}$
SIREN–2 $8.07\times 10^{-5}$ $7.09\times 10^{-3}$ $3.43\times 10^{-2}$
V Relu $2.45\times 10^{-1}$ $3.97\times 10^{-1}$ $9.99\times 10^{-1}$
V Tan $2.64\times 10^{-4}$ $9.95\times 10^{-3}$ $6.56\times 10^{-2}$
(c) Generalized Black–Scholes PDE
@@
(d) Burgers’ PDE
Models MSE MAE Max Error
FLM $6.55\times 10^{-3}$ $4.07\times 10^{-2}$ $3.73\times 10^{-1}$
SIREN–1 $2.13\times 10^{-2}$ $1.15\times 10^{-1}$ $3.89\times 10^{-1}$
SIREN–2 $9.80\times 10^{-3}$ $5.86\times 10^{-2}$ $3.59\times 10^{-1}$
SIREN–3 $6.96\times 10^{-3}$ $4.34\times 10^{-2}$ $3.68\times 10^{-1}$
V Relu $3.09\times 10^{-2}$ $1.46\times 10^{-1}$ $3.74\times 10^{-1}$
V Tan $4.58\times 10^{-3}$ $2.88\times 10^{-2}$ $4.44\times 10^{-1}$
Figure 5: Solution surfaces of the Heat (a), Poisson (b), GBS (c), and Burgers’ (d) equations, showing exact solution (left), FLM approximation (middle), and absolute error (right).
@@
The numerical experiments demonstrate that FLMs consistently outperform other NN models across a set of multidimensional PDEs. While SIREN and vanilla feedforward NNs perform well when the target function contains similar mathematical behavior as the activation function they use, FLMs achieve significantly lower error metrics and faster convergence (see tables in Appendix C) regardless of the PDE. In terms of the MSE and MAE metrics, FLMs were shown to perform better than all models considered for each PDE, with the exception of Burgers’ PDE. This difficulty in approximating the solution of Burgers’ equation was anticipated, mainly because of the presence of a discontinuity. However, the FLM was still able to attain an MSE of the same order of magnitude of the best performing model, the Vanilla Tan NN. In terms of Max Error, FLMs also showed an overall superior performance, even though not always dominating the other models. In general, out of all of the competing models, SIREN NNs are the models with the closest performance to that of FLMs. In the next section, we apply the proposed FLM model to the solution of a family of OCPs.
5.2 Numerical Experiments for Optimal Control Problems
@@
Both NN and Fourier series have been proven to be very effective to solve OCPs in recent years (Effati and Pakdaman, 2013; Nicolosi et al., 2022; Nicolosi and Griffin, 2023; Nicolosi et al., 2025; Rubel and Nicolosi, 2025). To show the effectiveness of FLMs to solve OCPs, we consider the optimal control of a system governed by the dynamics of a 3–strategy odd–circulant evolutionary game, namely, a Rock–Paper–Scissors game (Griffin and Fan, 2022). The state of the system, a vector $\bm{u}(t)=(u_{1},u_{2},u_{3})^{\top}$ with $u_{1}+u_{2}+u_{3}=1$, represents the proportion of a population playing each strategy. The goal is to push these proportions toward the desired equilibrium point $\bm{u}_{eq}=(1/3,1/3,1/3)^{\top}$ representing an equally balanced population.
@@
The system’s evolution is described by the replicator dynamics and is influenced by an external control input, $\gamma(t)$. This control actively alters the game’s payoff matrix, which is defined as $\bm{A}_{3}(\gamma)=\bm{L}_{3}+\gamma\bm{M}_{3}$, where $\bm{L}_{3}$ is the standard payoff matrix and $\bm{M}_{3}$ is an actuating matrix and given by
$\bm{L}_{3}=\begin{pmatrix}0&-1&1\\ 1&0&-1\\ -1&1&0\end{pmatrix},\quad\bm{M}_{3}=\begin{pmatrix}0&0&1\\ 1&0&0\\ 0&1&0\end{pmatrix}$.
@@
The objective is to find an optimal control strategy $\gamma(t)$ that minimizes a cost functional over a time horizon $[0,T]$. This cost has two components: a term that penalizes the squared distance of the state $\bm{u}(t)$ from the target equilibrium $\bm{u}_{eq}$, and a term that penalizes the squared control effort $\gamma(t)^{2}$.
@@
An important point to note is that the given OCP possesses a cyclic symmetry. If $\bm{u}^{*}(t)=(u^{*}_{1}(t),u^{*}_{2}(t),u^{*}_{3}(t))^{\top}$ is the optimal state trajectory, $\gamma^{*}(t)$ is the optimal control and $J^{*}$ is the optimal objective value for an initial condition $\bm{u_{0}}=(u_{01},u_{02},u_{03})^{\top}$, then any cyclic permutation of this initial state will yield the same optimal control solution and objective value. Consequently, the new optimal state trajectory will be the corresponding cyclic permutation of the original trajectory, $\bm{u}^{*}(t)$. For instance, if the initial state is permuted to $(u_{02},u_{03},u_{01})$, the optimal trajectory becomes $(u^{*}_{2}(t),u^{*}_{3}(t),u^{*}_{1}(t))$, while the optimal control and objective value remain unchanged. The numerical approximation of the solution of the OCP with our proposed model is presented in the next example.
@@
Consider a class of the OCP described above, where the initial condition $\bm{u_{0}}$ varies withing a given range, written as
min $J(\bm{u},\gamma)=\int_{0}^{T}\left(\frac{1}{2}\|\bm{u}-\bm{u}_{eq}\|_{2}^{2}+\frac{r}{2}\gamma^{2}\right)dt$
s.t. $\dot{\bm{u}}=\bm{F}(\bm{u})+\gamma(t)\bm{G}(\bm{u})$ (16)
$\bm{u}(0)=\bm{u}_{0}\in U_{0},$
where $\bm{u}(t)\in\Delta_{2}$ lies on the 2–simplex. The vector fields, for $i=1,2,3,$ are given by
$\bm{F}_{i}(\bm{u})$ $=u_{i}\left((\bm{e}_{i}-\bm{u})^{T}\bm{L}_{3}\bm{u}\right),$ (17)
$\bm{G}_{i}(\bm{u})$ $=u_{i}\left((\bm{e}_{i}-\bm{u})^{T}\bm{M}_{3}\bm{u}\right).$
@@
Let $\widehat{u}_{i}(t,\bm{u_{0}};\Theta_{u_{i}})$ be the FLM that learns the approximation of state variable $u_{i}(t,u_{0})$, i.e., $\bm{u}\approx\bm{\widehat{u}}=\left(\widehat{u}_{1},\widehat{u}_{2},\widehat{u}_{3}\right)^{\top}$ and $\widehat{\gamma}(t,\bm{u_{0}};\Theta_{\gamma})$ be the FLM that learns the control variable $\gamma(t,u_{0})$ for any trajectory departing from $U_{0}$, where $\Theta=\{\Theta_{u_{1}},\Theta_{u_{2}},\Theta_{u_{3}},\Theta_{\gamma}\}$ is the collection of all network parameters of the respective networks. Given the complex nature of the OCP above, the training process is different than it was for the PDEs and so is the construction of the training loss function. We train all the FLM simultaneously by minimizing the training loss using a penalty method (Ch. 9 in Bazaraa et al. (2006)). We define an unconstrained loss function that incorporates the objective functional and penalties for dynamics and initial condition violations as
$\mathcal{L}(\Theta)=\frac{1}{N_{\text{IC}}}\sum_{\bm{u}_{0}\in U_{0}}\left[\mathcal{J}+\frac{1}{2}\bm{\mu^{\top}_{1}}\bm{\mathcal{V}}_{\text{dyn}}+\frac{1}{2}\bm{\mu^{\top}_{2}}\bm{\mathcal{V}}_{\text{init}}\right],$ (18)
@@
$\mathcal{J}(\Theta)$ $=\int_{0}^{T}\left(\frac{1}{2}\|\bm{\widehat{u}}-\bm{u}_{eq}\|^{2}+\frac{r}{2}\widehat{\gamma}^{2}\right)dt,$
$\bm{\mathcal{V}}_{\text{dyn}}(\Theta)$ $=\int_{0}^{T}\left(\dot{\widehat{\bm{u}}}-\bm{F}(\widehat{\bm{u}})-\widehat{\gamma}\bm{G}(\widehat{\bm{u}})\right)\odot\left(\dot{\widehat{\bm{u}}}-\bm{F}(\widehat{\bm{u}})-\widehat{\gamma}\bm{G}(\widehat{\bm{u}})\right)dt,$
$\bm{\mathcal{V}}_{\text{init}}(\Theta)$ $=\left(\bm{\widehat{u}}(0)-\bm{u}_{0}\right)\odot\left(\bm{\widehat{u}}(0)-\bm{u}_{0}\right),$
where $\bm{\mu_{1}},\bm{\mu_{2}}$ are vectors of positive penalty coefficients. Minimizing the total loss $\mathcal{L}$ with respect to $\Theta_{u}$ and $\Theta_{\gamma}$ forces the FLM to approximate the optimal state trajectory and control signal that solve the original OCP. Next, we consider two cases: a fixed initial condition case, for a given $\bm{u_{0}}$, and a for a range of initial conditions $U_{0}\subset\Delta_{2}$.
@@ 5.2.1 Fixed Initial Condition OCP
First we solve the OCP for fixed initial condition with a one–dimensional FLM, which is a NN with one hidden–layer and cosinecosine activation function. We consider three such FLMs, two for the state variables $u_{1}$ and $u_{2}$ and one for the control $\gamma$. The third state variable is calculated using the relation $u_{3}=1-u_{1}-u_{2}$. The training is performed with default learning rate $0.001$ and betas $\beta_{1}=0.9$ and $\beta_{2}=0.999$ with the ADAM optimizer. To test the accuracy of our model, we compare its performance against a numerical solution obtained via PMP to formulate a two–point boundary value problem, which was then solved numerically to a high precision (tolerance of $10^{-8}$) using the solve_bvp function from Python’s SciPy library. The FLM approximated solutions and the PMP–based numerical solutions for three different, fixed initial conditions are shown in Fig. 6. The approximated optimal objective function values have absolute percentage errors of $0.47\%$, $0.37\%$, and $0.34\%$ respectively. Fig. 7 shows the state trajectory for the initial point $\bm{u}_{0}=(0.20,0.20,0.60)^{\top}$. The parameters of the OCP, as in Griffin and Fan (2022), are chosen as $T=6.0$, $r=0.2$.